On May 1, 2021, Meta Computer, Inc., enters into a contract to sell 5,600 units of Comfort Office Keyboard to one of its clients, Bionics, Inc., at a fixed price of $95,200, to be settled by a cash payment on May 1. Delivery is scheduled for June 1, 2021. As part of the contract, the seller offers a 25% discount coupon to Bionics for any purchases in the next six months. The seller will continue to offer a 5% discount on all sales during the same time period, which will be available to all customers. Based on experience, Meta Computer estimates a 50% probability that Bionics will redeem the 25% discount voucher, and that the coupon will be applied to $56,000 of purchases. The stand-alone selling price for the Comfort Office Keyboard is $19.00 per unit. Required: 1. How many performance obligations are in this contract? 2. Prepare the journal entry that Meta would record on May 1, 2021. 3. Assume the same facts and circumstances as above, except that Meta gives a 5% discount option to Bionics instead of 25%. In this case, what journal entry would Meta record on May 1, 2021?
Answer #1

1. Two performance obligations:

i) Keyboard

ii) Customer option for a future discount.

2. Journal Entry

Transaction No Account Title and Explanation Debit Credit
Cash $ 95,200
Deferred Revenue - Key Board $90,440
Deferred Revenue - Discount Coupon $4,760

Calculation;

Keyboards = 5,600 * $19 = $ 106,400

Option = $56,000 * (.25 - .05 ) * .50 = $ 5,600

Allocation ;

Key boards = ( 106,400 / 106,400 + 5,600)

= 106,400 / 112,000 = .95 * 95,200

= $90,440

Option = (5,600 / 106,400 + 5,600)

   =5,600/ 112,000   

   = .05 * 95,200 = 4,760

3. Journal Entry

No Account Title and Explanation Debit Credit
1 Cash $95,200
Deferred Revenue Keyboard $95,200